OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 2x + 2y = –6 3x – 2y = 11

OpenStudy (anonymous):

all im looking for is the solution to the system?

OpenStudy (anonymous):

y=3-1x y=-11/2 + 3/2

OpenStudy (anonymous):

umm does that even make any sense?

OpenStudy (anonymous):

the answers suppose to be in the form of (x,y)

OpenStudy (anonymous):

Do you have a graphing calculator?

OpenStudy (anonymous):

If you do, graph both lines. The intersect point is the answer

Directrix (directrix):

(1 , -4) Graphing Technology (2x + 2y = –6) and (3x-2y=11) http://www.wolframalpha.com/input/?i=%282x+%2B+2y+%3D+%E2%80%936%29+and+%283x-2y%3D11%29

OpenStudy (anonymous):

thnks